can someone please check my answers
x=y to 2nd z to 3rd
if y=3 and z=2
MY ANSWER X=3*3*2*2*2=72; x=72
x=y to the 2nd*z over t
if y=2, z=8 and t=4
MY ANSWER X=2*2*8/4=8;X=8
yx to the 2nd=zt
if y=3, z=1 and t=27
MY ANSWER 3*3*3=1*27;x=3
The last problem has another solution: x=-3
All your work looks good.
Cheers,
Stan H

Answer by stanbon(75887)

x=y^2 z^3
if y=3 and z=2
MY ANSWER X=3*3*2*2*2=72; x=72
x=y^2 z/t
if y=2, z=8 and t=4
MY ANSWER X=2*2*8/4=8;X=8
yx^2 =zt
if y=3, z=1 and t=27
MY ANSWER 3*3*3=1*27;x=3
The last problem has another solution: x=-3
All your work looks good.
Cheers,
Stan H